a body A is projected vertically upward from the top of the tower with the velocity 40m/s, the tower being 180m high.After t seconds , Another body B is allowed to fall from the same point.both the bodies rrach the ground simultaneously.l . calculate time and velocity of A and B after reaching the ground?

Grade:12th pass

1 Answers

Siddhartha Tiwari
6 years ago
First calculate the time taken by B to reach the ground. The equation comes out to be t^2 =2*180/10. This implies that t=6s.Now write the equation of motion of A from start to end. Total time is t+6 seconds. So,The equation now looks like ........-180 = 40*(t+6) - (1/2) * 10 * (t+6)^2 Now solve this equation to get the value of t.It is coming out to be 5.21s. Now you can easily calculate the velocities of both the bodies at the end of the trajectories.